I bought a truck, paying $2,000 as a down payment. I still owe 80% of the selling price. What was the selling price of the truck
Answer by lenny460(1073)

I bought a truck, paying $2,000 as a down payment. I still owe 80% of the selling price. What was the selling price of the truck
Therefore $2,000 = 20% of the selling price.
The balance owed is 80%.
10% = 2,000/2 = $1,000.
1,000 * 10 = 10,000.
The selling price is $10,000.